Solve by substitution method
y=40+0.25x
y=35+0.30x

Found 2 solutions by ilovezvezda, Blued0517:
Answer by ilovezvezda(1)

Solve by substitution method
y=40+0.25x
y=35+0.30x
from the first equation, y-40/.25=x
now substitute x in the 2nd equation with this value to find y
y=35+.30(y-40/.25) (I'll now write .25 as 1/4 in the denominator)
y=35+.30(4y-160) (to get rid of the decimal, we'll multiply both sides with 10)
10y=350+3(4y-160)
10y=350+12y-480
group y's in one side
2y=130
y=65
from this, plug in the value of y in any of the equations to find the value of x.

Answer by Blued0517(1)